# POGEMA: A Benchmark Platform for Cooperative Multi-Agent Navigation

### v1

@@ 3.2 POGEMA Toolbox
The POGEMA Toolbox is a comprehensive framework designed to facilitate the testing of learning-based approaches within the POGEMA environment. This toolbox offers a unified interface that enables the seamless execution of any learnable MAPF algorithm in POGEMA. Firstly, the toolbox provides robust management tools for custom maps, allowing users to register and utilize these maps effectively within POGEMA. Secondly, it enables the concurrent execution of multiple testing instances across various algorithms in a distributed manner, leveraging Dask [^7] https://github.com/dask/dask for scalable processing. The results from these instances are then aggregated for analysis. Lastly, the toolbox includes visualization capabilities, offering a convenient method to graphically represent aggregated results through detailed plots. This functionality enhances the interpretability of outcomes, facilitating a deeper understanding of algorithm performance.
POGEMA Toolbox offers a dedicated tool for map generation, allowing the creation of three distinct types of maps: random, mazes and warehouse maps. All generators facilitates map creation using adjustable parameters such as width, height, and obstacle density. Additionally, maze generator includes specific parameters for mazes such as the number of wall components and the length of walls. The maze generator was implemented based on the generator provided in  [10]. POGEMA Toolbox [^8] https://pypi.org/project/pogema-toolbox can be installed using PyPI, and licenced under Apache License 2.0.
@@ 3.3 Baselines
POGEMA integrates a variety of MARL, hybrid and planning-based algorithms with the environment. These algorithms, recently presented, demonstrate state-of-the-art performance in their respective fields. Table 2 highlights the differences between these approaches. Some, such as LaCAM and RHCR, are centralized search-based planners. Other approaches, such as SCRIMP and DCC, while decentralized, still require communication between agents to resolve potential collisions. Learnable modern approaches for LifeLong MAPF that do not utilize communication include Follower [46], MATS-LP [47], and Switchers [48] (Assistant Switcher, Learnable Switcher). All these approaches utilize independent PPO [43] as the training method.
The following modern MARL algorithms are included as baselines: MAMBA [12], QPLEX [60], IQL [53], VDN [52], and QMIX [37]. For environment preprocessing, we used the preprocessing scheme provided in the Follower approach, enhancing it with the anonymous targets of other agents’ local observations. We utilized the official implementation of MAMBA, as provided by its authors [^9] https://github.com/jbr-ai-labs/mamba, and employed PyMARL2 framework [^10] https://github.com/hijkzzz/pymarl2 for establishing MARL baselines.
@@
MAMBA [12]
QPLEX [60]
IQL [53]
VDN [52]
QMIX [37]
Follower [46]
MATS-LP [47]
Switcher [48]
SCRIMP [61]
DCC [29]
LaCAM [33] - - - - -
RHCR [25] - - - - -
@@ 4.1 Dataset
  • Mazes – maps that encouter prolonged corridors with 1-cell width that require high level of cooperation between the agent to accomplish the mission. These maps are procedurally generated.
  • Random – one of the most commonly used type of maps, as they are easy to generate and allow to avoid overfitting to some special structure of the map. POGEMA ontains an integrated random maps generator, that allows to control the density of the obstacles.
  • Warehouses – this type of maps are usually used in the papers related to LifeLong MAPF. While there is no narrow passages, high density of the agents might significantly reduce the overall throughput, especially when agents are badly distributed along the map. These maps are also can be procedurally generated.
  • MovingAI – a set of maps from the existing benchmark widely used in MAPF community. The contained maps have different sizes and structures. It can be used to show how the approach deals with single-agent pathfinding and also deals with the maps that have out-of-distribution structure.
  • MovingAI-tiles – a modified MovingAI set of maps. Due to the large size of the original maps, it’s hard to get high density of the agents on them. To get more crowded maps, we slice the original maps on 16 pieces with $64\times 64$ size.
@@
Start and goal locations are generated via random generators. They are generated with fixed seeds, thus can be reproduced. It’s guaranteed, that each agent has its own goal location and the path to it from its start location exists. Examples of the maps are presented in Figure 2.
@@ 4.2 Metrics
The existing works related to solving MAPF problems evaluates the performance by two major criteria – success rate and the primary performance indicators mentioned above: sum-of-costs, makespan, throughput. These are directly obtainable from POGEMA. While these metrics allow to evaluate the algorithms at some particular instance, it’s might be difficult to get a high-level conclusion about the performance of the algorithms. Thus, we want to introduce several high-level metrics that covers multiple different aspects:
@@
Out-of-Distribution – how well the algorithm works on out-of-distribution maps. This metric is computed in the same way as Performance, with the only difference that the approaches are evaluated on a set of maps, that were not used during training phase and have different structure of obstacles. For this purpose we utilize maps from MovingAI-tiles set of maps.
@@
Cooperation – how well the algorithm is able to resolve complex situations. To evaluate this metric we run the algorithm on Puzzles set of maps and compare the obtained results with best solutions that were obtained by classical MAPF/LMAPF solvers.
$Cooperation_{LMAPF}=throughput/throughput_{best}$ (7)
Congestion – how well the algorithm distributes the agents along the map and reduces redundant waits, collisions, etc. To evaluate this metric we compute the average density of the agents presented in the observations of each agent and compare it to the overall density of the agents on the map.
$Congestion=\frac{\sum_{i\in agents}agents\_density(obs_{i})/agents\_density(map)}{|agents|}$ (8)
Pathfinding – how well the algorithm works in case of presence of a single agent on a large map. This metric is tailored to determine the ability of the approach to effectively lead agents to their goal locations. For this purpose we run the approaches on large city maps from MovingAI benchmark sets. The obtained solution cost (in fact - length of the path) should be optimal.
$Pathfinding=\begin{cases}1\text{ if path is optimal}\\ 0\text{ otherwise}\end{cases}$ (9)
@@ 4.3 Experimental Results
We have evaluated all the supported baseline algorithms (12 in total) on both MAPF and LMAPF setups on all 6 datasets. The results of this evaluation are presented in Fig.3. The details about number of maps, number of agents, seeds, etc. are given in the supplementary material (as well as details on how these results can be reproduced).
@@
In both setups, i.e. MAPF and LMAPF, the best results in terms of cooperation, out-of-distribution and performance metrics were obtained by centralized planners, i.e. LaCAM and RHCR respectively.
For MAPF tasks, LaCAM outperforms all other approaches on all metrics except congestion. It is hypothesized that in this approach, the even distribution of agents across the environment is not crucial due to its centralized nature, which efficiently resolves complex conflicts. Specialized learnable MAPF approaches, i.e., DCC and SCRIMP, take second place, showing close performance but with different specifics. DCC shows better results on out-of-distribution tasks and pathfinding tasks than SCRIMP, which is better at managing congestion. Surprisingly, the results of SCRIMP are inferior on pathfinding tasks, suggesting a problem with this approach in single-agent tasks that do not require communication, which can be an out-of-distribution setup for this algorithm. MARL algorithms such as QPLEX, VDN, and QMIX underperform in comparison with other approaches, exhibiting a significant gap in the results, which can be attributed to the absence of additional techniques used in hybrid approaches, despite incorporating preprocessing techniques from the Follower approach. This could suggest that the MARL community lacks large-scale approaches and benchmarks for them. Predictably, IQL shows the poorest performance, highlighting the importance of centralized training for multi-agent pathfinding (MAPF) tasks that require high levels of cooperation. The best results out of all MARL approaches are shown by MAMBA, which has non-zero results on the performance metric and better results in terms of the cooperation metric. Its results are still much worse than those obtained by hybrid methods, and it is not able to solve any instances in the out-of-distribution dataset.
@@
For LMAPF tasks, the centralized approach, RHCR, dominates in cooperation, out-of-distribution, performance, and pathfinding metrics. However, it significantly lags behind other approaches in terms of congestion and scalability metrics. Out of the non-centralized approaches, the best results, depending on the metric, are shown by either MATS-LP or Follower. While MATS-LP is better in terms of cooperation and pathfinding, the latter is better in terms of congestion and scalability metrics. It’s also worth noting that MATS-LP requires much more runtime than other approaches as it runs MCTS for each of the agents at each step, which takes time (see Appendix F for more details). There are also two more hybrid approaches – ASwitcher and LSwitcher – which differ in the way they switch between planning-based and learning-based parts. One of the reasons for their mediocre results is a total lack of global information, i.e., the Switcher approach assumes that the agents have no information about the global map; thus, each of them needs to reconstruct the map based on their local observations. Again, MARL approaches underperform in these scenarios, with QMIX and QPLEX showing comparable results. QMIX performs better in cooperation and out-of-distribution metrics, while QPLEX excels in performance. As in the MAPF setup, among all the MARL approaches, the best results are demonstrated by MAMBA, which is slightly better than others in terms of cooperation and out-of-distribution metrics and much better in terms of the performance metric.
5 Conclusion and Limitations
@@
This paper presents POGEMA – a powerful suite of tools tailored for creating, assessing, and comparing methods and policies in multi-agent navigation problems. POGEMA encompasses a fast learning environment and a comprehensive evaluation toolbox suitable for pure MARL, hybrid, and search-based solvers. It includes a wide array of methods as baselines. The evaluation protocol described, along with a rich set of metrics, assists in assessing the generalization and scalability of all approaches. Visualization tools enable qualitative examination of algorithm performance. Integration with the well-known MARL API and map sets facilitates the benchmark’s expansion. Existing limitations are two-fold. First, a conceptual limitation is that communication between the agents is not currently disentangled in POGEMA environment. Second, the technical limitations include the lack of JAX support and integration with other well-known GPU parallelization tools.
@@ Appendix A Evaluation Setup Details
POGEMA benchmark contains $6$ different sets of maps and all baseline approaches were evaluated on them either on MAPF or on LMAPF instances. Regardless the type of instances, number of maps, seeds and agents were the same. Table 3 contains all information about these numbers. In total, this corresponds to 3,376 episodes for each scenario type. Note that there is no MaxSteps (LMAPF) value for MovingAI set of maps. This set of maps was used only for pathfinding meta-metric, thus all approaches were evaluated only on MAPF instances with a single agent.
Table 3: Details about the instances on different sets of maps.
@@
MovingAI [1] 8 256$\times$256 10 2048 -
MovingAI-tiles [64, 128, 192, 256] 128 64$\times$64 1 256 256
@@ Appendix B Results for MAPF Benchmark
In this section, we present the extended results of the MAPF benchmark analysis, highlighting the performance, out-of-distribution handling, scalability, cooperation, congestion, and pathfinding capabilities of various approaches. The experiments were conducted on different map types and sizes, employing metrics such as SoC, CSR, and makespan to evaluate effectiveness. Detailed visual and tabular data illustrate how centralized and learnable approaches perform under various conditions.
@@ B.2 Out-of-Distribution
Out-of-Distribution metric was calculated on MovingAI-tiles dataset, that consists of pieces of cities maps with $64\times 64$ size. Due to much larger size compared to Mazes and Random maps, the amount of agents was also significantly increased. The results are presented in Figure 5. Here again centralized search-based planner, i.e. LaCAM, demonstrates the best results both in terms of CSR and SoC. Among hybrid methods, DCC completely outperforms the others (including SCRIMP) on the out-of-distribution dataset. MARL approaches are unable to solve any instance even with 64 agents.
Figure 5: Performance of MAPF approaches on MovingAI-tiles maps. These results were utilized to compute Out-of-Distribution metric. The shaded area indicates $95\%$ confidence intervals.
@@ B.4 Cooperation
Algorithm CSR SoC
DCC $0.72$ p m 0.04 $96.33$ p m 9.97
IQL $0.27$ p m 0.04 $316.87$ p m 14.47
LaCAM $0.96$ p m 0.02 $36.29$ p m 7.26
MAMBA $0.39$ p m 0.04 $177.64$ p m 14.68
QMIX $0.53$ p m 0.05 $246.11$ p m 17.10
QPLEX $0.33$ p m 0.04 $250.12$ p m 14.80
SCRIMP $0.82$ p m 0.04 $104.31$ p m 13.73
VDN $0.11$ p m 0.03 $336.99$ p m 12.45
@@
B.5 Congestion
Congestion is one of the meta-metrics that estimates how well the agents are distributed along the map. This metric indirectly influences the performance of the approach, as well-distributed agents in case of highly crowded instances allows to reduce the amount of collisions and redundant wait-actions. To compute this metric we utilize the results obtained on Warehouse map with the highest evaluated amount of agents - $192$. In Table 5, we present results for a range of agents from $64$ to $192$ in increments of $32$, not just for $192$ agents. In contrast to other metrics, that are computed as a ratio to the best obtained results, Congestion metric is computed as the ratio to average density of all agents on the map to average density of the agents in agents’ local observations.
Table 5: Average agents density by number of agents across algorithms for Warehouse map. Please note, only column with $192$ agents was utilized to compute Congestion metric.
Algorithm 64 Agents 96 Agents 128 Agents 160 Agents 192 Agents
DCC $0.094$ p m 0.001 $0.132$ p m 0.001 $0.170$ p m 0.001 $0.207$ p m 0.001 $0.241$ p m 0.001
IQL $0.076$ p m 0.001 $0.114$ p m 0.001 $0.163$ p m 0.002 $0.244$ p m 0.002 $0.299$ p m 0.002
LaCAM $0.082$ p m 0.001 $0.116$ p m 0.001 $0.149$ p m 0.001 $0.179$ p m 0.001 $0.207$ p m 0.001
MAMBA $0.101$ p m 0.001 $0.183$ p m 0.001 $0.266$ p m 0.002 $0.335$ p m 0.002 $0.389$ p m 0.002
QMIX $0.073$ p m 0.001 $0.103$ p m 0.001 $0.130$ p m 0.001 $0.154$ p m 0.001 $0.179$ p m 0.001
QPLEX $0.077$ p m 0.001 $0.113$ p m 0.001 $0.146$ p m 0.001 $0.175$ p m 0.001 $0.205$ p m 0.001
SCRIMP $0.074$ p m 0.001 $0.104$ p m 0.001 $0.127$ p m 0.001 $0.148$ p m 0.001 $0.173$ p m 0.001
VDN $0.071$ p m 0.001 $0.101$ p m 0.001 $0.130$ p m 0.001 $0.158$ p m 0.001 $0.188$ p m 0.001
B.6 Pathfinding

### v2

@@ 3.2 POGEMA Toolbox
The POGEMA Toolbox is a comprehensive framework designed to facilitate the testing of learning-based approaches within the POGEMA environment. This toolbox offers a unified interface that enables the seamless execution of any learnable MAPF algorithm in POGEMA. Firstly, the toolbox provides robust management tools for custom maps, allowing users to register and utilize these maps effectively within POGEMA. Secondly, it enables the concurrent execution of multiple testing instances across various algorithms in a distributed manner, leveraging Dask [^6] https://github.com/dask/dask for scalable processing. The results from these instances are then aggregated for analysis. Lastly, the toolbox includes visualization capabilities, offering a convenient method to graphically represent the aggregated results through detailed plots. This functionality enhances the interpretability of outcomes, facilitating a deeper understanding of algorithm performance.
POGEMA Toolbox offers a dedicated tool for map generation, allowing the creation of three distinct types of maps: random, mazes and warehouse maps. All generators facilitates map creation using adjustable parameters such as width, height, and obstacle density. The maze generator was implemented based on the generator provided in  (Damani et al., 2021).
@@ 3.3 Baselines
POGEMA integrates a variety of MARL, hybrid and planning-based algorithms with the environment. These algorithms, recently presented, demonstrate state-of-the-art performance in their respective fields. Table 2 highlights the differences between these approaches. Some, such as LaCAM and RHCR, are centralized search-based planners. Other approaches, such as SCRIMP and DCC, while decentralized, still require communication between agents to resolve potential collisions. Learnable modern approaches for LifeLong MAPF that do not utilize communication include Follower (Skrynnik et al., 2024a), MATS-LP (Skrynnik et al., 2024b), and Switchers (Skrynnik et al., 2023). All these approaches utilize independent PPO (Schulman et al., 2017) as the training method.
MAMBA (Egorov & Shpilman, 2022)
QPLEX (Wang et al., 2020b)
IQL (Tan, 1993)
VDN (Sunehag et al., 2018)
QMIX (Rashid et al., 2020)
Follower (Skrynnik et al., 2024a)
MATS-LP (Skrynnik et al., 2024b)
Switcher (Skrynnik et al., 2023)
SCRIMP (Wang et al., 2023)
DCC (Ma et al., 2021)
LaCAM (Okumura, 2024; 2023) - - - - -
RHCR (Li et al., 2021b) - - - - -
@@
The following modern MARL algorithms are included as baselines: MAMBA (Egorov & Shpilman, 2022), QPLEX (Wang et al., 2020b), IQL (Tan, 1993), VDN (Sunehag et al., 2018), and QMIX (Rashid et al., 2020). For environment preprocessing, we used the scheme provided in the Follower approach, enhancing it with the anonymous targets of other agents’ local observations for MAPF scenarios. We utilized the official implementation of MAMBA, as provided by its authors [^7] https://github.com/jbr-ai-labs/mamba, and employed PyMARL2 framework [^8] https://github.com/hijkzzz/pymarl2 for establishing MARL baselines. We used the default parameters for MAMBA, since Dreamer (which serves as the foundation for MAMBA) is known to work effectively across domains with nearly the same hyperparameters. For the other MARL approaches, we performed a hyperparameter sweep over the learning rate, batch size, replay buffer size, and GRU hidden state size, using the best parameters based on the performance scores from the training scenarios on the Random and Mazes maps.
@@ 4.1 Dataset
  • Mazes – maps that encounter prolonged corridors with 1-cell width that require high level of cooperation between the agents. These maps are procedurally generated.
  • Random – one of the most commonly used type of maps, as they are easy to generate and allow to avoid overfitting to some special structure of the map. POGEMA contains an integrated random maps generator, that allows to control the density of the obstacles.
  • Warehouse – this type of maps are usually used in the papers related to LifeLong MAPF. While there is no narrow passages, high density of the agents might significantly reduce the overall throughput, especially when agents are badly distributed along the map.
  • Cities – a set of city maps from MovingAI – the existing benchmark widely used in heuristic-search community Sturtevant (2012). The contained maps have a varying structure and $256\times 256$ size. It can be used to show how the approach deals with single-agent pathfinding and also deals with the maps that have out-of-distribution structure.
  • Cities-tiles – a modified Cities set of maps. Due to the large size of the original maps, it’s hard to get high density of the agents on them. To get more crowded maps, we slice the original maps on 16 pieces with $64\times 64$ size.
@@
Start and goal locations are generated via random generators. They are generated with fixed seeds, thus can be reproduced. It is guaranteed that each agent has its own goal location and the path to it from its start location exists. Examples of the maps are presented in Figure 2.
@@ 4.2 Metrics
The existing works related to solving MAPF problems evaluates the performance by two major criteria – success rate and the primary performance indicators mentioned above: sum-of-costs, makespan, throughput. These are directly obtainable from POGEMA. While these metrics allow to evaluate the algorithms at some particular instance, it might be difficult to get a high-level conclusion about the performance. Thus, we want to introduce high-level metrics that cover multiple different aspects:
@@
Out-of-Distribution – how well the algorithm works on out-of-distribution maps. This metric is computed in the same way as Performance, with the only difference that the approaches are evaluated on a set of maps, that were not used during training phase and have different structure of obstacles. For this purpose we utilize maps from Cities-tiles set of maps.
@@
Cooperation – how well the algorithm is able to resolve complex situations. To evaluate this metric we run the algorithm on Puzzles set of maps.
$Cooperation_{LMAPF}=throughput/throughput_{best}$ (6)
@@
Coordination – how well the algorithm avoids collisions between agents and obstacles. Majority of learning-based approaches are decentralized, so there may be cases where multiple agents attempt to occupy the same cell or traverse the same edge simultaneously. There also might be the case when an agent tries to move to the blocked cell. The fewer collisions that occur during the episode, the better. To compute this metric, we use the results obtained on the Mazes and Random sets of maps.
$Coordination=1-\frac{occured\_collisions}{|agents|\times episode\_length}$ (8)
Pathfinding – how well the algorithm works in case of presence of a single agent on a large map. This metric is tailored to determine the ability of the approach to effectively lead agents to their goal locations. For this purpose we run the approaches on large maps from Cities benchmark sets. The closer the costs of the found paths to the optimal ones – the higher the score.
$Pathfinding=\begin{cases}path\_cost/path\_cost_{optimal}\\ 0\text{ if path not found}\end{cases}$ (9)
First three metrics, i.e. Performance, Out-of-Distribution, and Cooperation, have the same formula but differ in the set of maps used to compute them, while the remaining three metrics, i.e. Scalability, Coordination, and Pathfinding, are tailored to specific aspects.
@@ 4.3 Experimental Results
We have evaluated all the supported baselines (12 in total) on both MAPF and LMAPF setups on all 6 datasets. The results of this evaluation are presented in Figure 3. The details about number of maps, number of agents, seeds, etc. are given in the supplementary material (as well as details on how our results can be reproduced). In both setups, i.e., MAPF and LMAPF, the best results in terms of cooperation, out-of-distribution, and performance metrics were obtained by centralized planners, i.e. LaCAM and RHCR respectively.
@@
For MAPF tasks, LaCAM notably outperforms all other approaches. Specialized learnable MAPF approaches, such as DCC and SCRIMP, follow closely, with similar overall performance but different strengths. DCC performs better than SCRIMP on out-of-distribution tasks and pathfinding tasks, while the rest of the results are fairly similar. Surprisingly, SCRIMP underperforms on pathfinding tasks, indicating a weakness in single-agent tasks that do not require communication, which may be an out-of-distribution scenario for this algorithm. It is also worth noting that SCRIMP includes an integrated tie-breaking mechanism that ensures collision-free actions. MARL algorithms such as QPLEX, VDN, and QMIX perform significantly worse than other approaches, showing a substantial gap in the results. This can be attributed to the absence of additional techniques used in hybrid approaches. This may suggest that the MARL community lacks large-scale approaches and benchmarks for these tasks. Among MARL approaches, MAMBA achieves the best results in performance, cooperation, and pathfinding metrics, which can be attributed to its communication mechanism. However, its performance remains much worse than that of specialized methods, and it fails to solve any instances in the out-of-distribution dataset.
@@
For LMAPF tasks, the centralized approach, RHCR, is superior in all cases except for the scalability metric. Among the non-centralized approaches, the best results, depending on the metric, are shown by either MATS-LP or Follower. The only metric in which these approaches differ significantly is the cooperation metric, where MATS-LP performs better than Follower. However, MATS-LP requires considerably more test-time computation, as it runs MCTS for each agent at every step (see Appendix I for more details). Additionally, there are two hybrid approaches – ASwitcher and LSwitcher – that differ in how they alternate between planning-based and learning-based components. One reason for their mediocre performance is the lack of global information, i.e., they assume that agents have no prior knowledge of the map, requiring each agent to reconstruct it based on the local observations. Differently from the MAPF scenarios, MARL approaches can compete with hybrid methods on LMAPF instances. This behavior is partly explained by the use of Follower’s observation model, which is specifically designed to solve LMAPF. Among MARL approaches, QPLEX delivers the best results, in contrast to MAPF tasks it even outperforming MAMBA.
5 Conclusion, Limitations and Future Work
@@
This paper presents POGEMA – a powerful suite of tools tailored for creating, assessing, and comparing methods and policies for multi-agent pathfinding. POGEMA encompasses a fast learning environment and a comprehensive evaluation toolbox suitable for pure MARL, hybrid, and search-based solvers. It includes a wide array of methods as baselines. The evaluation protocol described, along with a rich set of metrics, assists in assessing the generalization and scalability of all approaches. Visualization tools enable qualitative examination of algorithm performance. Integration with the well-known MARL API and map sets facilitates the benchmark’s expansion. Existing limitations are two-fold. First, a conceptual limitation is that communication between agents is not currently disentangled in POGEMA environment. Second, technical limitations include the lack of JAX support and integration with the other GPU parallelization tools.
Future work could explore large-scale training setups for MARL methods, capable of handling large agent populations and scenarios, particularly within the CTDE (Centralized Training, Decentralized Execution) paradigm. Second, advancing communication learning suited to large-scale settings, where local interactions are crucial, is another promising direction. Third, addressing memory limitations by developing efficient approaches for long-horizon tasks without relying on global guidance is also vital. Furthermore, leveraging POGEMA’s procedural map generation and expert data from centralized solvers could enhance imitation learning and aid in training decentralized foundation models for MAPF. Finally, studying heterogeneous policy coordination may enable effective collaboration among diverse agents and advance concurrent training of different algorithms in shared environments.
6 Acknowledgments
We thank the researchers using POGEMA in their studies, as well as the repository users who contributed issues, questions, and feedback to improve the platform. Our thanks also go to the anonymous reviewers for their valuable comments and suggestions, which enhanced the quality of this work.
@@ 6 Acknowledgments
Appendix J
Ingestion of MovingAI Maps
Appendix L
POGEMA Speed Performance Evaluation
@@ Appendix A Evaluation Setup Details
POGEMA benchmark contains $6$ different sets of maps and all baseline approaches were evaluated on them either on MAPF or on LMAPF instances. Regardless the type of instances, number of maps, seeds and agents were the same. Table 3 contains all information about these numbers. In total, this corresponds to 3,376 episodes for each scenario type. Note that there is no MaxSteps (LMAPF) value for Cities set of maps. This set of maps was used only for pathfinding meta-metric, thus all approaches were evaluated only on MAPF instances with a single agent.
We used implementation LaCAM-v3999https://github.com/Kei18/lacam3, RHCR [^10] https://github.com/Jiaoyang-Li/RHCR. For learning-based approaches beyond MARL, we used their official implementations and the provided weights for Follower [^11] https://github.com/AIRI-Institute/learn-to-follow, MATS-LP [^12] https://github.com/CognitiveAISystems/mats-lp, Switcher [^13] https://github.com/AIRI-Institute/when-to-switch, SCRIMP [^14] https://github.com/marmotlab/SCRIMP, and DCC [^15] https://github.com/ZiyuanMa/DCC.
@@
Cities [1] 8 256$\times$256 10 2048 -
Cities-tiles [64, 128, 192, 256] 128 64$\times$64 1 256 256
Table 3: Details about the instances on different sets of maps.
@@ Appendix B Results for MAPF Benchmark
In this section, we present the extended results of the MAPF benchmark analysis, highlighting the performance, out-of-distribution handling, scalability, cooperation, coordination, and pathfinding capabilities of various approaches. The experiments were conducted on different map types and sizes, employing metrics such as SoC, CSR, and makespan to evaluate effectiveness. Detailed visual and tabular data illustrate how centralized and learnable approaches perform under various conditions.
@@ B.2 Out-of-Distribution
Out-of-Distribution metric was calculated on Cities-tiles dataset, that consists of pieces of cities maps with $64\times 64$ size. Due to much larger size compared to Mazes and Random maps, the amount of agents was also significantly increased. The results are presented in Figure 5. Here again centralized search-based planner, i.e. LaCAM, demonstrates the best results both in terms of CSR and SoC. Among hybrid methods, DCC completely outperforms the others (including SCRIMP) on the out-of-distribution dataset. MARL approaches are unable to solve any instance even with 64 agents.
Figure 5: Performance of MAPF approaches on Cities-tiles maps. These results were utilized to compute Out-of-Distribution metric. The shaded area indicates $95\%$ confidence intervals.
@@ B.4 Cooperation
Algorithm CSR SoC
DCC $0.75$ p m 0.04 $89.36$ p m 10.34
IQL $0.47$ p m 0.02 $250.03$ p m 18.83
LaCAM $1.00$ p m 0.00 $20.84$ p m 1.75
MAMBA $0.40$ p m 0.02 $173.00$ p m 13.92
QMIX $0.35$ p m 0.04 $253.28$ p m 16.24
QPLEX $0.51$ p m 0.05 $234.79$ p m 18.89
SCRIMP $0.85$ p m 0.03 $93.18$ p m 13.23
VDN $0.45$ p m 0.05 $242.42$ p m 16.87
@@
B.5 Pathfinding
Algorithm Makespan
DCC $189.56$ p m 26.29
IQL $1096.86$ p m 196.97
LaCAM $179.82$ p m 20.97
MAMBA $416.45$ p m 139.34
QMIX $1055.75$ p m 193.03
QPLEX $795.09$ p m 187.72
SCRIMP $1460.04$ p m 180.43
VDN $1114.21$ p m 211.93